OpenStudy (anonymous):

What is the EXACT area of the circle having a radius of 11 miles?

OpenStudy (anonymous):

A = 380.13

OpenStudy (ipwnbunnies):

^^ Not exactly exact. A = pi*radius^2

OpenStudy (johnweldon1993):

Right...so the EXACT area (since pi is not exact!) would be \[\large A = \pi r^2\] \[\large A = 11^2 \pi\] \[\large A = 121 \pi\]

OpenStudy (anonymous):

Is it 379.94

OpenStudy (ipwnbunnies):

No, it's 121pi. That is exact, since pi is transcendental, you can't express it with a few numbers.